OpenStudy (unklerhaukus):

please help \[{1 \over π}\int\limits_{-π}^{π}|\sin(x)|\cos(nx) dx\]

OpenStudy (unklerhaukus):

ok so the integrand is an even function right

OpenStudy (anonymous):

doing fourier series?

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

\[{1 \over π} ∫_0^π 2sin(x)cos(nx)dx\]

OpenStudy (anonymous):

you are trying to find \[a_n\] right?

OpenStudy (unklerhaukus):

and using a trig formulae.. 2sin(x)cos(nx) = sin((1+n)x)+sin((1-n)x)

OpenStudy (anonymous):

doing fourier series?

OpenStudy (anonymous):

I think it should be \[{1 \over π}\int\limits_{-π}^{π}|\sin( n x)|\cos(nx) dx\] because \[{1 \over π}\int\limits_{-π}^{π}|\sin(x)|\cos(nx) dx\]=0 by orthagonality

OpenStudy (unklerhaukus):

yeah i am trying to find the Fourier series of the absolute value of the sine wave i already have \[a_0 = 4/π\] and \[b_n = 0\] even function

OpenStudy (anonymous):

actually , you might not need to do all that since fourier series just respresentation using trig

OpenStudy (unklerhaukus):